find surface area of a cylinder circumference is 6in height is 8in
Answer by ewatrrr(24785)


Hi
C = 2pir r = 6/2pi = 3/pi
Surface Area = 2*base area + lateral area(C*h)
SA = 2pi*(3/pi)^2 + 6*8 = 18/pi + 48